OpenStudy (anonymous):

Melanie is thinking of a four-digit number. The digit in the thousands place is double the digit in the hundreds place. The digit in the hundreds place is double the digits in the tens place. The digit in the tens place is double the digit in the ones place. There is a 1 in the ones place. What is the number?

OpenStudy (anonymous):

8421

OpenStudy (anonymous):

Reverse engineering is how you do it.

OpenStudy (anonymous):

thk you mathsolver ur lifesaver

OpenStudy (anonymous):

say the number is (abcd) where a is the digit in the thousands place,b is the hudreds place, c is the tens place and d is the ones place. according to you d=1,c=2*d=2,b=2*c=2*2=4,a=2*b=2*4=8,,,,,,hence the number is 8421...hope this helps